Required information [The following information applies to the questions displayed below.] Solomon Company makes and sells prReq B1 Req B2 Prepare an income statement using the contribution margin format. answers to nearest whole number.) SOLOMON COMc. Suppose that Solomon desires to earn a $312,000 profit. Determine the sales volume in units and dollars required to earn tReq c1 Req c2 Suppose that Solomon desires to earn a $312,000 profit. Prepare an income statement using the contribution marge. If fixed costs drop to $310,000, what level of sales is required to earn the desired profit? Express your answer in unitse. If fixed costs drop to $310,000, what level of sales is required to earn the desired profit? Express your answer in unitsg. Assume that Solomon concludes that it can sell 13,000 units of product for $80 each. Recall that variable costs are $30 eaPlease help me. I thought I was doing good but then idk what happened.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Solomon
Answer b 1 Amount $ Note
Sell price                    102.00 A
Less: Variable costs                       24.00 B
Contribution margin per unit                       78.00 C
Total Fixed costs             413,400.00 D
Break even units                 5,300.00 E=D/C
Break even $            540,600.00 F=E*A
Answer b 2 Amount $
Sales             540,600.00 See F
Less: Variable costs             127,200.00 G=B*E
Contribution margin            413,400.00 H=F-G
Total Fixed costs             413,400.00 See D
Net Income (Loss)                             -   I=H-D
Answer c 1 Amount $
Total Fixed costs             413,400.00 See D
Add: Desired Profit             312,000.00
Desired Contribution margin            725,400.00 J
Contribution margin per unit                       78.00 See C
Sales units                 9,300.00 K=J/C
Sales $            948,600.00 L=K*A
Answer c 2 Amount $
Sales             948,600.00 See L
Less: Variable costs             223,200.00 M=K*B
Contribution margin            725,400.00
Less: Total Fixed costs             413,400.00
Net Income (Loss)            312,000.00
Answer e 1 Amount $
Revised Fixed costs             310,000.00
Add: Desired Profit             312,000.00
Desired Contribution margin            622,000.00 N
Contribution margin per unit                       78.00 See C
Sales units                 7,975.00 O=N/C
Sales $            813,450.00 P=O*A
Answer e 2 Amount $
Sales             813,450.00 See P
Less: Variable costs             191,400.00 Q=O*B
Contribution margin            622,050.00
Less: Revised Fixed costs             310,000.00
Net Income (Loss)            312,050.00
Answer g Amount $
Total units sold               13,000.00 R
Less: Break even units                 5,300.00 See E
Margin of safety                 7,700.00 S=R-E
Margin of safety $            785,400.00 T=S*A
Margin of safety % 59% U=S/R